# Heavy baryons in the relativized quark model with chromodynamics

### v1

@@ Abstract
Following the work of Capstick and Isgur [Phys. Rev. D 34, 2809 (1986)], we systematically study the mass spectrum of the heavy baryons in the relativized quark potential model with chromodynamics. Besides the original Godfrey-Isgur (GI) model, we also adopt a modified GI model which replaces the linear confinement by a screened one. The two models give similar results in our work. All heavy baryons observed so far can be explained as three-quark states. In particular, we identify the $\Omega_{c}(3000)$/$\Omega_{b}(6316)$, $\Omega_{c}(3050)$/$\Omega_{b}(6330)$, $\Omega_{c}(3065)$/$\Omega_{b}(6340)$ and $\Omega_{c}(3090)$/$\Omega_{b}(6350)$ states as the $p_{\lambda}$ excitations with quantum numbers $1/2^{-}$, $3/2^{-}$, $3/2^{-}$ and $5/2^{-}$. The $\Omega_{c}(3120)$ is a $3/2^{-}$ state with the $p_{\rho}$ excitation, whose bottom partner is predicted to be $\Omega_{b}(6446/6457,3/2^{-})$. The higher state $\Omega_{c}(3188)$ is the $2s_{\lambda}$ excitation with quantum numbers $1/2^{+}$, and $\Omega_{c}(3327)$ is a $d_{\lambda}$ excitation with quantum numbers $3/2^{+}$ or $5/2^{+}$. In addition, the $\Lambda_{c}(2940)$ with quantum numbers $J^{P}=3/2^{-}$ could be explained as the $p_{\rho}$ excitation.
@@ I Introduction
In the past few decades, we have seen large experimental progress in searching for the heavy baryons. Lots of charmed baryons were observed in the last few years. In 2017, the LHCb collaboration observed five narrow states in the $\Xi_{c}^{+}K^{-}$ mass spectrum: $\Omega_{c}(3000)^{0}$, $\Omega_{c}(3050)^{0}$, $\Omega_{c}(3066)^{0}$, $\Omega_{c}(3090)^{0}$ and $\Omega_{c}(3119)^{0}$ [1], where the first four of them were confirmed by the Belle collaboration [2]. Later LHCb observed $\Lambda_{c}(2860)^{+}$ with quantum numbers $J^{P}=3/2^{+}$ in the $D^{0}p$ channel [3]. In 2020, LHCb further studied the $\Lambda_{c}^{+}K^{-}$ mass spectrum and observed $\Xi_{c}(2923)^{0}$, $\Xi_{c}(2939)^{0}$ and $\Xi_{c}(2965)^{0}$ [4], where the first two are new states. In 2022, the Belle collaboration observed $\Lambda_{c}(2910)^{+}$ decaying into to $\Sigma_{c}\pi$ [5]. Recently, the LHCb collaboration observed two new states $\Omega_{c}(3185)^{0}$ and $\Omega_{c}(3227)^{0}$ in the $\Xi_{c}^{+}K^{-}$ decay modes [6]. In addition, many bottom baryons were also found in experiments. In 2018, the LHCb collaboration reported the observation of $\Sigma_{b}(6097)^{\pm}$ in the $\Lambda_{b}^{0}\pi^{\pm}$ final states [7]. The same year, they found the $\Xi_{b}(6227)^{-}$ as a peak in both the $\Lambda_{b}^{0}K^{-}$ and $\Xi_{b}^{0}\pi^{-}$ invariant mass spectra [8]. The isospin partner $\Xi_{b}(6227)^{0}$ was soon found [9]. In 2019, the LHCb collaboration observed the $\Lambda_{b}(6146)^{0}$ and $\Lambda_{b}(6152)^{0}$ in the $\Lambda_{b}^{0}\pi^{+}\pi^{-}$ channel [10]. Another state $\Lambda_{b}(6072)^{0}$ was also found in the same channel by the CMS collaboration[11]. In 2020, the LHCb collaboration observed four narrow peak, $\Omega_{b}(6316)^{-}$, $\Omega_{b}(6330)^{-}$, $\Omega_{b}(6340)^{-}$ and $\Omega_{b}(6350)^{-}$, in the $\Xi_{b}^{0}K^{-}$ channel [12]. Later in 2021, the CMS collaboration found $\Xi_{b}(6100)^{-}$ in the $\Xi_{b}^{-}\pi^{+}\pi^{-}$ channel [13]. Recently, the LHCb collaboration confirmed the existence of $\Xi_{b}(6227)^{0}$ and observed another state $\Xi_{b}(6333)^{0}$ [14]. To date, there are 42 singly-charmed and 25 singly-bottomed baryons collected in PDG [15], plus 3 singly-charmed and 1 singly-bottomed baryons observed more recently.
Inspired by the experimental progress, physicists used various methods to study heavy baryons, such as the quark models [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], lattice QCD [39, 40, 41, 42, 43, 44, 45], QCD sum rules [46, 47, 48, 49, 50, 51, 52], heavy quark effective theory (HQET) [53], and the Regge phenomenology [54, 55, 56]. For extra references, see recent review Refs. [57, 58, 59, 60, 61, 62].
@@ II.1 Godfrey-Isgur (GI) model
The $V_{\text{string}}$ interaction can be broken into an effective two-body piece and a three-body piece:
@@ II.3 Baryon wave function
The baryons should be color-singlet, thus we have only one totally anti-symmetric color wave function
@@ III.1 Parameters
The GI model is successful in studying the mesons and baryons. In 1985, Godfrey and Isgur fitted the parameters for mesons [63]. Later Capstick and Isgur re-fitted the parameters for the baryons [70]. The concrete values of these parameters are collected in Table 1. After their works, lots of heavy baryons had been observed in experiments [15]. To appreciate the experimental progress in the past three decades, we perform the $\chi^{2}$ fits to obtain the 13/14 parameters of the GI/MGI models. The $\chi^{2}$ is defined as
@@ III.2 The $\Sigma_{Q}$ and $\Lambda_{Q}$ baryons
First we consider the $\Sigma_{c}$ baryons. From Fig. 3, we can clearly see that the low-lying states form band structures. The lowest band consists of the ground states of spin $1/2^{+}$ and $3/2^{+}$. The next band consists of seven negative parity states dominated by $P$-wave excitation components. Among them, the $1/2^{-}$ and $3/2^{-}$ states further form a two semi-band structure. The lower two states are dominated by $p_{\lambda}$ modes, while the higher ones are dominated by $p_{\rho}$ excitations. In 2004, the Belle Collaboration observed the $\Sigma_{c}(2800)$ state in the $\Lambda_{c}\pi$ channel [92]. Its quantum numbers are not determined yet [15]. Our calculation suggests that it is most likely a $3/2^{-}$ state with the $p_{\lambda}$ excitation. The third band is dominated by the $N=2$ excitation (For convenience, we use the oscillator band to label the quark model states. When we refer to a state within $N=2$ band, we are implying that most part of its wave functions belong to the $N=2$ band of oscillator). More precisely, this band consists of $2s_{\rho}$, $2s_{\lambda}$, $d_{\rho}$, $d_{\lambda}$ and $[p_{\rho}p_{\lambda}]_{L}$ excitations. We again find that the excitation over $\bm{\lambda}$ mode is easier than over $\bm{\rho}$ mode. Taking $1/2^{+}$ states as an example, the $\Sigma_{c}(2925/2923)$ and $\Sigma_{c}(3025/3032)$ are dominated by the $2s_{\lambda}$ and $d_{\lambda}$ modes, the $\Sigma_{c}(3096/3112)$ and $\Sigma_{c}(3119/3118)$ are dominated by $d_{\rho}$ and $2s_{\rho}$ modes, while the $\Sigma_{c}(3025/3040)$ and $\Sigma_{c}(3165/3180)$ are both dominated by the $[p_{\rho}p_{\lambda}]_{L}$ mode. We also calculate the heavier states which include higher excitations. They are more packed with each other. A more detailed study of their decay properties is needed.
@@ III.3 The $\Omega_{Q}$ baryons
For the $\Omega_{c}$ system (see Fig. 9), the $\Omega_{c}$ and $\Omega_{c}^{*}$ are ground states, which are used for fits in this work. Next we consider the $P$-wave excitation with negative parity, namely the seven states lying in the $3.0\sim 3.15~{}\mathrm{GeV}$ mass region. Among them, the $\Omega_{c}(3097/3103,1/2^{-})$ and $\Omega_{c}(3132/3141,3/2^{-})$ states are $p_{\rho}$ excitations, while the others are $p_{\lambda}$ one. As shown in Fig. 9, these states may correspond to the five narrow $\Omega_{c}$ excited states observed by the LHCb collaboration [1]. Note the the $\Omega_{c}(3020/3025,1/2^{-})$ and $\Omega_{c}(3033/3041,1/2^{-})$ states are dominated by the mixing of
@@
bases in the notation |$\ket{(l_{\rho},l_{\lambda})_{L},S_{12},j_{q}=L+S_{12},J^{P}}$. The calculations based on ${}^{3}P_{0}$ model suggested that a pure $\psi_{1}$ state is very broad [$\Gamma\sim\mathcal{O}(10^{2}~{}\text{MeV})$] while a pure $\psi_{2}$ state is very narrow ($\Gamma\sim 0~{}\text{MeV}$[96]. Since our parameters are solely fitted from the masses, we are unable to resolve the mixing angle between these two states with small mass difference. We can only say that one of the two states is broad. More precisely, the $\Omega_{c}(3000)$ is a $p_{\lambda}$ excitation with quantum numbers $1/2^{-}$, though we can not determine whether it corresponds to $\Omega_{c}(3020/3025,1/2^{-})$ or $\Omega_{c}(3033/3041,1/2^{-})$. The $\Omega_{c}(3050)$ and $\Omega_{c}(3065)$ correspond to the $\Omega_{c}(3054/3059,3/2^{-})$ and $\Omega_{c}(3073/3080,3/2^{-})$, which are also $p_{\lambda}$ excitations. The $\Omega_{c}(3090)$ is close to $\Omega_{c}(3097/3103,1/2^{-})$ and $\Omega_{c}(3104/3110,5/2^{-})$ states. However, Ref. [96] suggested that the $J^{P}=1/2^{-}$ $p_{\rho}$ excitation should be a broad state ($\sim 10^{2}~{}\text{MeV}$). Thus we interpret $\Omega_{c}(3090)$ as the $p_{\lambda}$ excitation $\Omega_{c}(3104/3110,5/2^{-})$. The next state, $\Omega_{c}(3120)$, is the $p_{\rho}$ excitation $\Omega_{c}(3132/3143,3/2^{-})$. In the higher mass region, the LHCb collaboration found a broad structure around 3188 MeV [1], which was later confirmed by the Belle [2] and LHCb [6] collaborations. In Ref. [6], the LHCb collaboration also found the $\Omega_{c}(3327)$. Comparing their masses with our calculated masses, we interpret the $\Omega_{c}(3188)$ as the $2s_{\lambda}$ excitation $\Omega_{c}(3176/3178,1/2^{+})$, while $\Omega_{c}(3327)$ is a $d_{\lambda}$ excitation with quantum numbers $3/2^{+}$ or $5/2^{+}$.
@@
Now we consider the $\Omega_{b}$ system (see Fig. 10). First we consider the ground states. Only the $1/2^{+}$ state is observed in experiment, while the $3/2^{-}$ state, $\Omega_{b}^{*}$, is not observed yet. Our calculation predicts the mass of $\Omega_{b}^{*}$ to be 6078/6077 MeV. Next we consider the $P$-wave excitations. Unlike the $\Omega_{c}$ case, an explicit two semi-band structure appears among the $P$-wave $\Omega_{b}$ excited states. The five lower states are $p_{\lambda}$ excitations while the two higher states are $p_{\rho}$ excitations. In 2020, the LHCb collaboration observed four narrow excited $\Omega_{b}$ states, namely the $\Omega_{b}(6316)$, $\Omega_{b}(6330)$, $\Omega_{b}(6340)$ and $\Omega_{b}(6350)$ states [12]. As shown in Fig. 10, the masses of these states lie in the mass region of the $p_{\lambda}$ excitations. Similar to the logic of the $\Omega_{c}$ baryons, the $\Omega_{b}(6316)$ is a $1/2^{-}$ state. The $\Omega_{b}(6330)$ and $\Omega_{b}(6340)$ correspond to $\Omega_{b}(6341/6347,3/2^{-})$ and $\Omega_{b}(6355/6360,3/2^{-})$. And the $\Omega_{b}(6350)$ corresponds to $\Omega_{b}(6369/6374,5/2^{-})$. The $p_{\rho}$ excitations is about 100 MeV above the $p_{\lambda}$ excitations. In particular, the $\Omega_{b}(6446/6457,3/2^{-})$, which is the bottom counterpart of the $\Omega_{c}(3120)$ state, should be a narrow state. We hope the future experiments can search for these higher states.

### v2

@@ Abstract
Following the work of Capstick and Isgur [Phys. Rev. D 34, 2809 (1986)], we systematically study the mass spectrum of the heavy baryons in the relativized quark potential model with chromodynamics. Besides the original Godfrey-Isgur (GI) model, we also adopt a modified GI model which replaces the linear confinement by a screened one. The two models give similar results in our work. All heavy baryons observed so far can be explained as three-quark states. In particular, we identify the $\Omega_{c}(3000)$/$\Omega_{b}(6316)$, $\Omega_{c}(3050)$/$\Omega_{b}(6330)$, $\Omega_{c}(3065)$/$\Omega_{b}(6340)$, and $\Omega_{c}(3090)$/$\Omega_{b}(6350)$ states as the $p_{\lambda}$ excitations with quantum numbers $1/2^{-}$, $3/2^{-}$, $3/2^{-}$, and $5/2^{-}$. The $\Omega_{c}(3120)$ is a $3/2^{-}$ state with the $p_{\rho}$ excitation, whose bottom partner is predicted to be $\Omega_{b}(6446/6457,3/2^{-})$. The higher state $\Omega_{c}(3188)$ is the $2s_{\lambda}$ excitation with quantum numbers $1/2^{+}$, and $\Omega_{c}(3327)$ is a $d_{\lambda}$ excitation with quantum numbers $3/2^{+}$ or $5/2^{+}$.
@@ I Introduction
In the past few decades, we have seen large experimental progress in searching for the heavy baryons. Lots of charmed baryons were observed in the last few years. In 2017, the LHCb collaboration observed five narrow states in the $\Xi_{c}^{+}K^{-}$ mass spectrum: $\Omega_{c}(3000)^{0}$, $\Omega_{c}(3050)^{0}$, $\Omega_{c}(3066)^{0}$, $\Omega_{c}(3090)^{0}$, and $\Omega_{c}(3119)^{0}$ [1], where the first four of them were confirmed by the Belle collaboration [2]. Later LHCb observed $\Lambda_{c}(2860)^{+}$ with quantum numbers $J^{P}=3/2^{+}$ in the $D^{0}p$ channel [3]. In 2020, LHCb further studied the $\Lambda_{c}^{+}K^{-}$ mass spectrum and observed $\Xi_{c}(2923)^{0}$, $\Xi_{c}(2939)^{0}$ and $\Xi_{c}(2965)^{0}$ [4], where the first two are new states. In 2022, the Belle collaboration observed $\Lambda_{c}(2910)^{+}$ decaying into $\Sigma_{c}\pi$ [5]. Recently, the LHCb collaboration observed two new states $\Omega_{c}(3185)^{0}$ and $\Omega_{c}(3227)^{0}$ in the $\Xi_{c}^{+}K^{-}$ decay modes [6]. In addition, many bottom baryons were also found in experiments. In 2018, the LHCb collaboration reported the observation of $\Sigma_{b}(6097)^{\pm}$ in the $\Lambda_{b}^{0}\pi^{\pm}$ final states [7]. The same year, they found the $\Xi_{b}(6227)^{-}$ as a peak in both the $\Lambda_{b}^{0}K^{-}$ and $\Xi_{b}^{0}\pi^{-}$ invariant mass spectra [8]. The isospin partner $\Xi_{b}(6227)^{0}$ was soon found [9]. In 2019, the LHCb collaboration observed the $\Lambda_{b}(6146)^{0}$ and $\Lambda_{b}(6152)^{0}$ in the $\Lambda_{b}^{0}\pi^{+}\pi^{-}$ channel [10]. Another state $\Lambda_{b}(6072)^{0}$ was also found in the same channel by the CMS collaboration[11]. In 2020, the LHCb collaboration observed four narrow peak, $\Omega_{b}(6316)^{-}$, $\Omega_{b}(6330)^{-}$, $\Omega_{b}(6340)^{-}$, and $\Omega_{b}(6350)^{-}$, in the $\Xi_{b}^{0}K^{-}$ channel [12]. Later in 2021, the CMS collaboration found $\Xi_{b}(6100)^{-}$ in the $\Xi_{b}^{-}\pi^{+}\pi^{-}$ channel [13]. Recently, the LHCb collaboration confirmed the existence of $\Xi_{b}(6227)^{0}$ and observed another state $\Xi_{b}(6333)^{0}$ [14]. To date, there are 42 singly-charmed and 25 singly-bottomed baryons collected in PDG [15], plus 3 singly-charmed and 1 singly-bottomed baryons observed more recently.
Inspired by the experimental progress, physicists used various methods to study heavy baryons, such as the quark models [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], lattice QCD [42, 43, 44, 45, 46, 47, 48], QCD sum rules [49, 50, 51, 52, 53, 54, 55], heavy quark effective theory (HQET) [56], and the Regge phenomenology [57, 58, 59]. For extra references, see recent review Refs. [60, 61, 62, 63, 64, 65].
@@ II.1 Godfrey-Isgur (GI) model
The $V_{\text{string}}$ interaction can be broken into an effective two-body piece and a three-body piece
@@ II.3 Baryon wave function
The baryons should be color-singlet, thus we have only one totally antisymmetric color wave function
@@ III.1 Parameters
The GI model is successful in studying the mesons and baryons. In 1985, Godfrey and Isgur fitted the parameters for mesons [66]. Later Capstick and Isgur refitted the parameters for the baryons [73]. The concrete values of these parameters are collected in Table 1. After their works, lots of heavy baryons had been observed in experiments [15]. To appreciate the experimental progress in the past three decades, we perform the $\chi^{2}$ fits to obtain the 13/14 parameters of the GI/MGI models. The $\chi^{2}$ is defined as
@@ III.2 The $\Sigma_{Q}$ and $\Lambda_{Q}$ baryons
First we consider the $\Sigma_{c}$ baryons. From Fig. 3, we can clearly see that the low-lying states form band structures. The lowest band consists of the ground states of spin $1/2^{+}$ and $3/2^{+}$. The next band consists of seven negative parity states dominated by $P$-wave excitation components. Among them, the $1/2^{-}$ and $3/2^{-}$ states further form a two semi-band structure. The lower two states are dominated by $p_{\lambda}$ modes, while the higher ones are dominated by $p_{\rho}$ excitations. In 2004, the Belle Collaboration observed the $\Sigma_{c}(2800)$ state in the $\Lambda_{c}\pi$ channel [95]. Its quantum numbers are not determined yet [15]. Our calculation suggests that it is most likely a $3/2^{-}$ state dominated by the $p_{\lambda}$ excitation. The third band is dominated by the $N=2$ excitation (For convenience, we use the oscillator band to label the quark model states. When we refer to a state within $N=2$ band, we are implying that most part of its wave functions belong to the $N=2$ band of oscillator). More precisely, this band consists of $2s_{\rho}$, $2s_{\lambda}$, $d_{\rho}$, $d_{\lambda}$ and $[p_{\rho}p_{\lambda}]_{L}$ excitations. We again find that the excitation over ${\lambda}$ mode is easier than over ${\rho}$ mode. Taking $1/2^{+}$ states as an example, the $\Sigma_{c}(2925/2923)$ and $\Sigma_{c}(3025/3032)$ are dominated by the $2s_{\lambda}$ and $d_{\lambda}$ modes, the $\Sigma_{c}(3096/3112)$ and $\Sigma_{c}(3119/3118)$ are dominated by $d_{\rho}$ and $2s_{\rho}$ modes, while the $\Sigma_{c}(3025/3040)$ and $\Sigma_{c}(3165/3180)$ are both dominated by the $[p_{\rho}p_{\lambda}]_{L}$ mode. We also calculate the heavier states which include higher excitations. They are more packed with each other. A more detailed study of their decay properties is needed.
@@ III.3 The $\Omega_{Q}$ baryons
For the $\Omega_{c}$ system (see Fig. 9), the $\Omega_{c}$ and $\Omega_{c}^{*}$ are ground states, which are used for fits in this work. Next we consider the $P$-wave excitation with negative parity, namely the seven states lying in the $3.0\sim 3.15~{}\mathrm{GeV}$ mass region. Among them, the $\Omega_{c}(3097/3103,1/2^{-})$ and $\Omega_{c}(3132/3141,3/2^{-})$ states are dominated by the $p_{\rho}$ excitations, while the others are $p_{\lambda}$ one. As shown in Fig. 9, these states may correspond to the five narrow $\Omega_{c}$ excited states observed by the LHCb collaboration [1]. Note that the $\Omega_{c}(3020/3025,1/2^{-})$ and $\Omega_{c}(3033/3041,1/2^{-})$ states are dominated by the mixing of
@@
bases in the notation |$\ket{(l_{\rho},l_{\lambda})_{L},S_{12},j_{q}=L+S_{12},J^{P}}$. The calculations based on ${}^{3}P_{0}$ model suggested that a pure $\psi_{1}$ state is very broad [$\Gamma\sim\mathcal{O}(10^{2}~{}\text{MeV})$] while a pure $\psi_{2}$ state is very narrow ($\Gamma\sim 0~{}\text{MeV}$[99]. Since our parameters are solely fitted from the masses, we are unable to resolve the mixing angle between these two states with small mass difference. We can only say that one of the two states is broad. More precisely, the $\Omega_{c}(3000)$ is a $1/2^{-}$ state dominated by the $p_{\lambda}$ excitation, though we cannot determine whether it corresponds to $\Omega_{c}(3020/3025,1/2^{-})$ or $\Omega_{c}(3033/3041,1/2^{-})$. The $\Omega_{c}(3050)$ and $\Omega_{c}(3065)$ correspond to the $\Omega_{c}(3054/3059,3/2^{-})$ and $\Omega_{c}(3073/3080,3/2^{-})$, which are also dominated by the $p_{\lambda}$ excitations. The $\Omega_{c}(3090)$ is close to $\Omega_{c}(3097/3103,1/2^{-})$ and $\Omega_{c}(3104/3110,5/2^{-})$ states. However, Ref. [99] suggested that the $J^{P}=1/2^{-}$ $p_{\rho}$ excitation should be a broad state ($\sim 10^{2}~{}\text{MeV}$). Thus we interpret $\Omega_{c}(3090)$ as the $\Omega_{c}(3104/3110,5/2^{-})$ state, which is dominated by the $p_{\lambda}$ excitation . The next state, $\Omega_{c}(3120)$, corresponds to the $p_{\rho}$-dominated excitation $\Omega_{c}(3132/3143,3/2^{-})$. In the higher mass region, the LHCb collaboration found a broad structure around 3188 MeV [1], which was later confirmed by the Belle [2] and LHCb [6] collaborations. In Ref. [6], the LHCb collaboration also found the $\Omega_{c}(3327)$. Comparing their masses with our calculated masses, we interpret the $\Omega_{c}(3188)$ as the $\Omega_{c}(3176/3178,1/2^{+})$ state dominated by the $2s_{\lambda}$ excitation, while $\Omega_{c}(3327)$ is a $d_{\lambda}$-dominated excitation with quantum numbers $3/2^{+}$ or $5/2^{+}$.
@@
Now we consider the $\Omega_{b}$ system (see Fig. 10). First we consider the ground states. Only the $1/2^{+}$ state is observed in experiment, while the $3/2^{-}$ state, $\Omega_{b}^{*}$, is not observed yet. Our calculation predicts the mass of $\Omega_{b}^{*}$ to be 6078/6077 MeV. Next we consider the $P$-wave excitations. Unlike the $\Omega_{c}$ case, an explicit two semi-band structure appears among the $P$-wave $\Omega_{b}$ excited states. The five lower states are dominated by $p_{\lambda}$ excitations while the two higher states are dominated by $p_{\rho}$ excitations. In 2020, the LHCb collaboration observed four narrow excited $\Omega_{b}$ states, namely the $\Omega_{b}(6316)$, $\Omega_{b}(6330)$, $\Omega_{b}(6340)$, and $\Omega_{b}(6350)$ states [12]. As shown in Fig. 10, the masses of these states lie in the mass region of the $p_{\lambda}$-dominated excitations. Similar to the logic of the $\Omega_{c}$ baryons, the $\Omega_{b}(6316)$ is a $1/2^{-}$ state. The $\Omega_{b}(6330)$ and $\Omega_{b}(6340)$ correspond to $\Omega_{b}(6341/6347,3/2^{-})$ and $\Omega_{b}(6355/6360,3/2^{-})$. And the $\Omega_{b}(6350)$ corresponds to $\Omega_{b}(6369/6374,5/2^{-})$. The $p_{\rho}$-dominated excitations is about 100 MeV above the $p_{\lambda}$-dominated excitations. In particular, the $\Omega_{b}(6446/6457,3/2^{-})$, which is the bottom counterpart of the $\Omega_{c}(3120)$ state, should be a narrow state. We hope the future experiments can search for these higher states.